Write point slope form of an equation of a line for
(2,-6) (2,5)

Answer by lwsshak3(11628)

Write point slope form of an equation of a line for
(2,-6) (2,5)
***
point slope form of equation for a line: y=mx+b, m=slope, b=y-intercept
slope=∆y/∆x=(5-(-6))/(2-2)=11/0=undefined
So, there is no point slope form of equation
if you draw a line through the 2 points, you will get a vertical line, x=2
Equation for given line: x=2